JET-COOLED NO_{2} SPECTRUM AROUND THE DISSOCIATION THRESHOLD D_{0} (∼ 25128 cm^{−1})

SYLVAIN HEILLIETTE, ANTOINE DELON, RÉMY JOST and \underline{PATRICK DUPRÉ}, Grenoble High Magnetic Field Laboratory, CNRS-MPI, BP 166, 25 Rue des Martyrs 38042 GRENOBLE, Cedex 9, France.

The \underline{C}avity \underline{R}ing \underline{D}own \underline{S}pectroscopy (CRDS) technique is a unique tool for probing non fluorescing absorbing molecular species. High resolution in achievable in the blue-UV energy region by intracavity frequency doubling a CW laser source (here a Ti:Sa) and by using a seeded supersonic jet (slit nozzle) expansion. The NO_{2} radical is known to strongly absorb in the blue energy range. However, if a LIF signal can be easily detected up to the dissociation threshold D_{0} (NO_{2} → NO(^{2}Π_{1/2}) + O(^{3}P_{2})), above, a lack of fluorescence is observed which is typical of a photodissociation process. The usual techniques for level detection above the dissociation threshold (PHOFEX or Fluorescence Depletion Pumping, for example) are based on pulsed sources (laser bandwidth limited). At the opposite, the CW CRDS technique (residual Doppler width: ∼ 400 MHz) allows to probe resonances above the threshold without laser bandwidth limitation. Resonances, whose the width spreads from ∼ 0.055 cm^{−1} (corresponding to a dissociation time of ∼ 200 ps) just above D_{0}, to larger shapes (∼ 1 cm^{−1}) without clear structure 10 cm^{−1} higher, are identified.